OpenStudy (anonymous):

Find the GCF of each pair of numbers a)28,84 b)65,39c)75,45

OpenStudy (wolf1728):

I guess you mean Greatest Common Factor? To do this, separate the numbers into their prime factors 28 = 2 * 2 * 7 84 = 2 * 2 * 3 * 7 We see the factors they have in common are 2 * 2 * 7 =28 So their greatest common factor is 28. Try this with b) and c)